# Retrieval-Augmented Review Generation for Poisoning Recommender Systems

### v1

@@ V-C1 Overall Imperceptibility
Since we are the first to incorporate textual reviews into fake user profiles to enhance attack effectiveness, existing detectors, which primarily focus on rating-only profiles (e.g., [19, 6]), are inadequate for identifying RAGAN. To demonstrate the imperceptibility of the attack, we employ t-SNE [6] to visualize the real user’s and RAGAN’s profile representations generated by the guardian module on real-world datasets, which involves rating and review information. As shown in Fig. 4, the representations of fake user profiles are scattered within the distribution of real user profiles on real-world datasets. The overall proximity between attack profiles and normal profiles in the embedding space can be attributed to the imperceptibility optimizations, which enforces high semantic similarity. Meanwhile, the dispersed nature of the attack profiles may stem from our style transfer strategy, which enhances the diversity and realism of the generated behaviors. Hence, RAGAN can effectively mimic real users, enabling highly inconspicuous attacks.
V-C2 Review Quality
To further assess the imperceptibility of the attack, we conduct an additional evaluation on the generation quality of the adversarial reviews. Following the works [53, 46], we adopt the text perplexity score computed by GPT-Neo as the metric. A lower perplexity indicates that the perturbed text remains closer to natural human language, thus appearing less suspicious and detectable. As shown in Fig. 5, RAGAN achieves a lower average perplexity than R-Trojan, demonstrating its superiority in generating high-quality review text. These reviews are not only fluent and coherent but also semantically meaningful and stealthy, thereby enhancing their deceptive effectiveness. The performance gap may be attributed to the stronger ICL and generative abilities of RAGAN’s underlying multimodal FMs powered by the retrieval demonstration algorithm and the text style transfer strategy, compared to R-Trojan’s plain GPT-2.
TABLE V: Ablation and parameter studies of RAGAN on real-world datasets.
@@ V-C2 Review Quality
Dataset Component WRMF NCF LightGCN DeepCoNN DeepCoNN++
$\lambda$ $p$ $\widetilde{p}_{k}$ HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
Amazon 0.5 1 $\times$ 0.5581 0.3137 0.3527 0.2078 0.1967 0.1079 0.4064 0.1191 0.6081 0.1901
0 1 0.4673 0.2619 0.2605 0.1316 0.1407 0.0741 0.3019 0.0882 0.4369 0.1316
0.5 0.5 0.5784 0.3650 0.3904 0.2271 0.2199 0.1160 0.5210 0.1772 0.7163 0.2291
0.5 1 0.6313 0.4254 0.4209 0.2526 0.2569 0.1331 0.5624 0.2018 0.7438 0.2334
Yelp 0.5 1 $\times$ 0.1859 0.1340 0.1593 0.0750 0.1023 0.0650 0.2267 0.0666 0.3303 0.1099
0 1 0.1276 0.0938 0.1250 0.0659 0.0654 0.0459 0.0848 0.0248 0.1587 0.0468
0.5 0.5 0.2383 0.1523 0.1865 0.1362 0.1159 0.0836 0.2383 0.0701 0.4268 0.1431
0.5 1 0.2494 0.2011 0.2040 0.1469 0.1308 0.0961 0.2597 0.0760 0.4495 0.1394
@@
TABLE VI: Examples of fake ratings with the corresponding fake textual reviews from different attacks on real-world datasets.
@@
V-D Ablation Studies
We remove some pivotal components of RAGAN and investigate the performance changes. Table V illustrates the performance of ablation studies of RAGAN. We give results on two representative datasets, which are Amazon Musical Instruments and Yelp, due to space limitations.
To validate the effectiveness of our review quality enhancement component, we conduct ablation experiments by removing the multimodal demonstration retrieval algorithm. Since our previous work [21] has empirically demonstrated the effectiveness of textual reviews for strengthening fake user profiles, we focus solely on assessing how review quality affects attack performance. By comparing the cases of $\widetilde{p}_{k}=\times$ and $\widetilde{p}_{k}=\checkmark$ (i.e., zero-shot vs. few-shot learning), we observe that improving the quality of fake textual reviews contributes notably to the transferability of the attack. The main reason may be that the reviews encapsulate rich item features, which can help focus on a broader set of users with similar taste preferences and thus enhance item exposure rates. Moreover, as shown in Table III, Table IV and Table V, RAGAN without $\widetilde{p}_{k}$ achieves comparable performance to R-Trojan across various victim RSs. This is primarily attributed to the use of powerful foundation models with extensive corpus, which possess strong capabilities in understanding, reasoning and generation, thereby producing more semantically meaningful review text than the GPT-2 backbone used in R-Trojan. Moreover, as shown in Table V, removing the instructional agent (i.e., $\lambda=0$) results in a significant drop in attack performance against different victim models, indicating its vital role in optimizing and enhancing the attack transferability.
V-E Attack Performance with Partial Knowledge
@@
To evaluate the effectiveness of RAGAN in more practical settings, we conduct experiments with the assumption that the attacker can access only 50% of the user-item interaction records, i.e., $p=0.5$. Due to space limitations, we focus on representative datasets, reporting mainly on Amazon Musical Instruments and additionally on Yelp for a more comprehensive evaluation. For the other datasets, the trends are similar. From Fig. 4(d) and Table V, we can observe that RAGAN is still effective against various black-box RSs on real-world datasets, and its profiles remain similar to normal user profiles, with high-quality reviews generated. These results indicate that RAGAN can generate transferable and imperceptible fake user profiles, even when only partial knowledge of the training data of the victim RSs is accessed. This weak model in Table V is significantly superior to all the baselines learned with the full data in Table III and Table IV, and the same advantage is also shown in Fig. 5, demonstrating the effectiveness of the semantically rich textual reviews in reinforcing the attack. These findings highlight that existing RSs are vulnerable to semantically informed threat models in realistic scenarios.
V-F Examples of Adversarial Attack Profiles
@@
We provide examples to demonstrate the quality of the attack profiles generated by R-Trojan and RAGAN (note that R-Trojan is the first and only method to incorporate reviews into the profiles in the literature). Due to space constraints, we randomly select one item from the profiles on each type of dataset. As illustrated in Table VI, the reviews in R-Trojan are incoherent, factually inaccurate, semantically confused, logically inconsistent, sparse in content and misaligned with the ratings. In contrast, the reviews in RAGAN are more semantically meaningful, richer in valuable features, more coherent and natural, and better aligned with the ratings, resulting in higher-quality profiles. This observation also sheds light on why RAGAN is effective and why its profiles are both transferable and imperceptible.
VI Conclusion
@@
Many existing attack methods are overly idealized and fail to reflect realistic adversarial scenarios, which can result in overly optimistic assessments of the robustness of RSs in practice. To mitigate this problem, in this paper, we propose a novel poisoning attack framework, RAGAN, for generating high-quality fake user profiles including carefully designed ratings and textual reviews. RAGAN achieves significant improvements over existing profiles, particularly in the coherence and semantic richness of the reviews, as well as in the alignment between ratings and reviews. As a result, RAGAN demonstrates state-of-the-art transferability and imperceptibility in a more practical setting, as shown in the comprehensive experiments on real-world datasets. RAGAN enables us to delve into the vulnerabilities of current RSs, revealing their inherent security risks under specific input features (e.g., well-crafted ratings and/or reviews), and guides us to secure them in actual deployment scenarios.

### v2

@@ 5.3.1 Overall Imperceptibility
Since we are very early attempts to incorporate textual reviews into fake user profiles to enhance attack effectiveness, existing detectors, which currently focus on rating-only profiles (e.g., [19, 6]), have semantic gaps for identifying RAGAN. To demonstrate the imperceptibility of the attack, we employ t-SNE [6], which is a typical method used in the current works [21, 48], to visualize the real user’s and RAGAN’s profile representations generated by the guardian module on real-world datasets, which involves rating and review information. As shown in Fig. 4, the representations of fake user profiles are scattered within the distribution of real user profiles on real-world datasets. The overall proximity between attack profiles and normal profiles in the embedding space can be attributed to the imperceptibility optimizations, which enforces high semantic similarity. Meanwhile, the dispersed nature of the attack profiles may stem from our style transfer strategy, which enhances the diversity and realism of the generated behaviors. Hence, RAGAN can effectively mimic real users, enabling highly inconspicuous attacks.
Table 5: Ablation studies of RAGAN on real-world datasets.
@@
Dataset Component WRMF NCF LightGCN DeepCoNN Review-based Agent
HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
Amazon RAGANa 0.5581 0.3137 0.3527 0.2078 0.1967 0.1079 0.4064 0.1191 0.6081 0.1901
RAGANb 0.4673 0.2619 0.2605 0.1316 0.1407 0.0741 0.3019 0.0882 0.4369 0.1316
RAGAN 0.6313 0.4254 0.4209 0.2526 0.2569 0.1331 0.5624 0.2018 0.7438 0.2334
Yelp RAGANa 0.1859 0.1340 0.1593 0.0750 0.1023 0.0650 0.2267 0.0666 0.3303 0.1099
RAGANb 0.1276 0.0938 0.1250 0.0659 0.0654 0.0459 0.0848 0.0248 0.1587 0.0468
RAGAN 0.2494 0.2011 0.2040 0.1469 0.1308 0.0961 0.2597 0.0760 0.4495 0.1394
@@
Table 6: Examples of fake ratings with the corresponding fake textual reviews from different attacks on real-world datasets.
@@
5.3.2 Review Quality
To further assess the imperceptibility of the attack, we conduct an additional evaluation on the generation quality of the adversarial reviews. Following the works [53, 46], we adopt the text perplexity score computed by GPT-Neo as the metric. A lower perplexity indicates that the perturbed text remains closer to natural human language, thus appearing less suspicious and undetectable. As shown in Fig. 5, RAGAN achieves a lower average perplexity than R-Trojan, demonstrating its superiority in generating high-quality review text. These reviews are not only fluent and coherent but also semantically meaningful and stealthy, thereby enhancing their deceptive effectiveness. This performance advantage arises because RAGAN effectively harnesses the full potential of its multimodal FMs powered by the retrieval demonstration algorithm and the text style transfer strategy, thereby achieving stronger ICL and generative capabilities than R-Trojan’s plain prompting approach.
5.4 Attack Performance with Partial Knowledge
@@
To evaluate the effectiveness of RAGAN in more practical settings, we conduct experiments with the assumption that the attacker can access only 50% of the user-item interaction records, i.e., $p=0.5$. Due to space limitations, we focus on representative datasets, reporting mainly on Amazon Musical Instruments and additionally on Yelp for a more comprehensive evaluation. For the other datasets, the trends are similar. From Fig. 6 and Fig. 4(d), we can observe that RAGAN is still effective against various black-box RSs on real-world datasets, and its profiles remain similar to normal user profiles, with high-quality reviews generated. These results indicate that RAGAN can generate transferable and imperceptible fake user profiles, even when only partial knowledge of the training data of the victim RSs is accessed. This weak model in Fig. 6 is significantly superior to all the baselines learned with the full data in Table 3 and Table 4, and the same advantage is also shown in Fig. 5, demonstrating the effectiveness of the semantically rich textual reviews in reinforcing the attack. These findings highlight that existing RSs are vulnerable to semantically informed threat models in realistic scenarios.
5.5 Ablation Studies
We remove some pivotal components of RAGAN and investigate the attack performance changes. Table 5 illustrates the performance of ablation studies of RAGAN. We give results on two representative datasets, which are Amazon Musical Instruments and Yelp, due to space limitations.
To validate the effectiveness of our review quality enhancement component, we conduct ablation experiments by removing the multimodal demonstration retrieval algorithm (denoted as RAGANa). Since the previous work [21] has empirically demonstrated the effectiveness of textual reviews for strengthening fake user profiles, we focus solely on assessing how review quality affects attack performance. By comparing the cases of RAGANa (zero-shot) and the full RAGAN (few-shot), we observe that improving the quality of fake textual reviews contributes notably to the transferability of the attack. The main reason may be that the reviews encapsulate rich item features, which can help focus on a broader set of users with similar taste preferences and thus enhance item exposure rates. Moreover, as shown in Table 3, Table 4 and Table 5, RAGANa achieves comparable performance to R-Trojan across various victim RSs. This is primarily attributed to the task-specific prompting techniques tailored on the powerful FMs, which enhances the models’ capabilities in understanding, reasoning and generation, thereby producing more semantically meaningful review text than the simple prompting method used in R-Trojan. Moreover, removing the instructional agent (denoted as RAGANb) results in a significant drop in attack performance against different victim models, as shown in Table 5, indicating its vital role in optimizing and improving the transferability of the attack.
We do not ablate the Guardian module, as removing it would break the GAN framework and make the adversarial optimization meaningless. Similarly, the style transfer strategy is kept unchanged, because it is a prompt-level cue designed to add linguistic diversity for more natural review expressions rather than a learnable module. Empirically, its removal yields negligible changes in attack effectiveness.
5.6 Examples of Adversarial Attack Profiles
@@
We provide examples to demonstrate the quality of the attack profiles generated by R-Trojan and RAGAN (note that R-Trojan is the first and only method to incorporate reviews into the profiles in the literature). Due to space constraints, we randomly select one item from the profiles on each type of dataset. As illustrated in Table 6, the reviews in R-Trojan are incoherent, factually inaccurate, semantically confused, logically inconsistent, sparse in content and misaligned with the ratings. In contrast, the reviews in RAGAN are more semantically meaningful, richer in valuable features, more coherent and natural, and better aligned with the ratings, resulting in higher-quality profiles. This observation also sheds light on why RAGAN is effective and why its profiles are both transferable and imperceptible.
6 Conclusion and Future Works
@@
Many existing attack methods are overly idealized and fail to reflect realistic adversarial scenarios, which can result in overly optimistic assessments of the robustness of RSs in practice. To mitigate this problem, in this paper, we propose a novel poisoning attack framework, RAGAN, for generating high-quality fake user profiles including carefully designed ratings and textual reviews. RAGAN achieves significant improvements over existing profiles, particularly in the coherence and semantic richness of the reviews, as well as in the alignment between ratings and reviews. As a result, RAGAN demonstrates state-of-the-art transferability and imperceptibility in a more practical setting, as shown in the comprehensive experiments on real-world datasets. RAGAN enables us to delve into the vulnerabilities of current RSs, revealing their inherent security risks under specific input features (e.g., well-crafted ratings and/or reviews), and guides us to secure them in actual deployment scenarios. Since our works are conducted on real-world datasets, future work will extend the RAGAN framework to real-world RSs for realistic robustness analysis and systematic defense testing, paving the way toward more secure practical deployments. Beyond this, we plan to 1) explore adaptive defense mechanisms that dynamically counteract evolving poisoning strategies, and 2) generalize RAGAN to multimodal and LLM-based recommenders, where textual, visual, and conversational signals jointly influence recommendations.